OpenStudy (anonymous):

Solve for y: y-3/y+7 - 5/y = 0 (fractions)

OpenStudy (anonymous):

\[\frac{y-3}{y+7}-\frac{5}{y}=0\] like that?

OpenStudy (anonymous):

yes!!

OpenStudy (saifoo.khan):

SATELLITE TO THE RESCUEE!

OpenStudy (anonymous):

wish I could write them like that!

OpenStudy (saifoo.khan):

WIRELESS rescue!

OpenStudy (anonymous):

get \[\frac{y-3}{y+7}=\frac{5}{y}\] so \[y(y-3)=5(y+7)\] \[y^2-3y=5y+35\] \[y^2-8y-35=0\]

OpenStudy (saifoo.khan):

solve for y, right.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes, solve for y

OpenStudy (saifoo.khan):

that isnt solve for y! ^ :(

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y=??

OpenStudy (saifoo.khan):

its like y = 11.14 y = -3.14

OpenStudy (anonymous):

i get \[y^2-8y=35\] so \[(y-4)^2=35+16=51\] \[y-4=\pm\sqrt{51}\] amd \[y=4\pm\sqrt{51}\]

OpenStudy (anonymous):

now let me check the algebra

OpenStudy (saifoo.khan):

^.^

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

Awesome! Except, lol, I'm not sure I understand the symbols in the answer... y=4/51 ???